OpenStudy (anonymous):

1. Joe is 5 years older than Sally. Represent Joe's age in terms of Sally's age. 2. The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. 3. Fifteen more than four times a number is 6 more than five times the number. 4. Martha and Sally had lunch together on Tuesday. Sally paid $2 more than Martha. The total bill was $5.64. 5. Bob paid $3 less than Tim at a local pizza parlor. Together they spent $7.80. 6. There were six more women on a committee than men. The committee was made up of 30 people. 7. Three consecutive integers have the sum of 153.

OpenStudy (anonymous):

1. sally age =x joe age=x+5

OpenStudy (anonymous):

there up i put the answers for these

OpenStudy (angela210793):

1- joe's age y=x+5 where x=sally's age 2- l=3+2w 3- 15+4x=6++5x 4- 2x+x=5.64 where X=amount in $ 5- (x-3)+x=7.80 6- x no. of men so : 6x+x=30 7- x+(x+1)+(x+2)=153 Now these r too easy I guess u can solve them by urself :)